### v2

# Blind Baselines Beat Membership Inference Attacks for Foundation Models

Blind Baselines Beat Membership Inference Attacks for Foundation Models
Debeshee Das, Jie Zhang, Florian Tramèr
ETH Zurich
Аннотация
Membership inference (MI) attacks try to determine if a data sample was used to train a machine learning model. For foundation models trained on unknown Web data, MI attacks are often used to detect copyrighted training materials, measure test set contamination, or audit machine unlearning. Unfortunately, we find that evaluations of MI attacks for foundation models are flawed, because they sample members and non-members from different distributions. For 8 published MI evaluation datasets, we show that blind attacks—that distinguish the member and non-member distributions without looking at any trained model—outperform state-of-the-art MI attacks. Existing evaluations thus tell us nothing about membership leakage of a foundation model’s training data.
1 Introduction
Many foundation models Bommasani et al. (2021) such as GPT-4 OpenAI (2023), Gemini Gemini Team (2023) and DALL-E Ramesh et al. (2021) are trained on unknown data. There is great interest in methods that can determine if a piece of data was used to train these models. Such methods—called membership inference attacks Shokri et al. (2017)—have been studied for foundation models in many recent works Shi et al. (2023); Duan et al. (2024); Ko et al. (2023); Dubiński et al. (2024); Zhang et al. (2024a); Meeus et al. (2023). Applications include privacy attacks Carlini et al. (2021), demonstrating the use of copyrighted material Meeus et al. (2024b), detecting test data contamination Oren et al. (2023), or auditing the efficacy of methods to ‘‘unlearn’’ training data Bourtoule et al. (2021).
To evaluate the performance of a membership inference attack, it is common to train a model on a random subset of a larger dataset, and then ask the attacker to distinguish these members from the remaining non-members by interacting with the model. Emulating this setup for foundation models is hard, since we often do not have access to a held out set sampled from the same distribution as the training set.
Existing MI evaluations thus create member and non-member datasets a posteriori, typically by picking members from sources known or suspected to be in the targeted foundation model’s training set, and then attempting to emulate the distribution of these sources to sample non-members.
Unfortunately, we show these strategies are severely flawed, and create easily distinguishable member and non-member distributions. As a special case of this flaw, concurrent works of Duan et al. (2024), Maini et al. (2024) and Meeus et al. (2024a) find a temporal shift between members and non-members in the Wiki-MIA dataset Shi et al. (2023). We show this issue is persistent by identifying significant distribution shifts (beyond temporal shifts) in 8 MI evaluation datasets for foundation models, for both text and vision. Worse, we show that existing MI attacks perform ‘‘worse than chance’’ on these datasets. Specifically, we design ‘‘blind’’ attacks, which completely ignore the target model, and outperform all reported results from state-of-the-art MI attacks (see Table 1).
Our methods are naive: for some datasets with a temporal shift, we apply a threshold to specific dates extracted from each sample; for other text or text-vision datasets, we build simple bag-of-words or n-gram classifiers on captions. We find that such simple methods work even for MI evaluation datasets that were explicitly designed to remove distribution shifts between members and non-members. Our work shows that removing biases in a post-hoc fashion is highly brittle.
Current MI attacks for foundation models thus cannot be relied on, as we cannot rule out that they are (poorly) inferring membership based on data features, without extracting any actual membership leakage from the model. In addition, studies that rely on the MI evaluations on these datasets, also cannot be trusted. For example, the authors of Panaitescu-Liess et al. (2024) investigate the effect of watermarking on MI attacks and evaluate it on datasets that we show to have clear distribution shifts. Future MI attacks should be evaluated on models with a clear train-test split, e.g., using the Pile Gao et al. (2020) or a random subset of DataComp Gadre et al. (2024) or DataComp-LM Li et al. (2024).
Таблица 1: MI evaluation datasets for foundation models are flawed. Our blind attacks distinguish members from non-members better than existing attacks, without looking at any model.
MI dataset Metric Best Reported (%) Ours (%)
WikiMIA TPR@5%FPR 43.2 94.7
BookMIA AUC ROC 88.0 91.4
Temporal Wiki AUC ROC 79.6 79.9
Temporal ArXiv AUC ROC 74.5 75.3
ArXiv (all vs 1 month) TPR@1%FPR 5.9 10.6
ArXiv (1 month vs 1 month) TPR@1%FPR 2.5 2.7
LAION-MI TPR@1%FPR 2.5 8.9
Gutenberg TPR@1%FPR 18.8 55.1
2 Background and Related Work
Web-scale training datasets.
Foundation models are often trained on massive datasets collected from web crawls, such as C4 Raffel et al. (2020), the Pile  Gao et al. (2020) or LAION Schuhmann et al. (2022). However, there has been a trend towards using undisclosed training sets for models like GPT-2 to GPT-4 Radford et al. (2019); OpenAI (2023), Gemini Gemini Team (2023) or DALL-E Ramesh et al. (2021). Even recent open models like LLaMA Touvron et al. (2023) do not release information about their training dataset. Some datasets have been released for research purposes, such as RedPajama Together (2023), Dolma Soldaini et al. (2024) or LAION Schuhmann et al. (2022). Notably, these datasets lack a designated test set.
Membership inference attacks.
Membership inference attacks aim to determine whether a given data point was used to train a machine learning model (Shokri et al., 2017). Early attacks applied a global decision function to all samples (e.g., by thresholding the model’s loss Yeom et al. (2018)). Current state-of-the-art attacks calibrate the attack threshold to the difficulty of each sample (Carlini et al., 2022).
Membership inference for foundation models.
Membership inference attacks have been applied to LLMs Shi et al. (2023); Duan et al. (2024); Zhang et al. (2024a); Meeus et al. (2023), diffusion models Dubiński et al. (2024), CLIP Ko et al. (2023), and other foundation models. The motivations for these attacks include using them as a component of a privacy attack Carlini et al. (2021), for evaluating unlearning methods Shi et al. (2023), or to detect the use of copyrighted data Meeus et al. (2024b). Due to the lack of a dedicated test set (and possibly even an unknown training set) for the targeted models, many of these works design custom evaluation datasets for membership inference attacks, by collecting sources of data that were likely used (respectively not used) for training.
Evaluating membership inference.
Membership inference attacks were originally evaluated with average-case metrics such as the ROC AUC on a balanced set of members and non-members Shokri et al. (2017). More recent work advocates for evaluating the attack’s performance in the worst-case, typically by reporting the true-positive rate (TPR) at low false-positive rates (FPR) Carlini et al. (2022).
Membership inference evaluations are usually set up so that a baseline attack (which does not query the target model) achieves 50% AUC (or equal TPR and FPR). Some authors have also considered cases where the attacker has a non-uniform prior Jayaraman et al. (2020). In either case, the goal of an MI attack is to extract information from the model to beat a baseline inference that does not have access to the model.
Many MI evaluation datasets for foundation models introduce distribution shifts between members and non-members, which allows for baseline attacks with non-trivial success. In concurrent work, Duan et al. (2024) and Maini et al. (2024) identify a temporal shift between members and non-members in one dataset—WikiMIA Shi et al. (2023)—and demonstrate that some attacks fail when temporal shifts are removed. Our work shows that this issue is much broader: virtually all evaluation sets proposed for membership inference on foundation models are flawed. We further show that existing attacks do not just exploit distribution shifts, but they do so sub-optimally and are easily beaten by blind baselines. These attacks thus perform ‘‘worse than chance’’.
3 Blindly Inferring Membership
3.1 Distribution Shifts in MI Evaluation Datasets
MI evaluations measure the ability to distinguish a model’s training set members from non-members. A baseline attack (with no knowledge of the model) should do no better than random guessing—i.e., an AUC of $50\%$ or equal TPR and FPR. For current MI evaluation datasets for foundation models, this is not the case, due to intrinsic differences between members and non-members. We discuss the most common reasons for this discrepancy below.
Temporal shifts.
MI evaluation sets based on a hard data cutoff for some evolving Web source (e.g., Wikipedia or arXiv) introduce a temporal shift between members and non-members. We can thus blindly distinguish members from non-members if we can answer a question of the form: ‘‘is this data from before or after 2023?’’
Biases in data replication.
Even if we know how the training set was sampled, building an indistinguishable dataset of non-members is challenging Recht et al. (2019). Slight variations in the procedures used to create the two datasets Engstrom et al. (2020) could be exploited by a blind attack.
Distinguishable tails.
Some works filter and process the data to maximally align the member and non-member distributions (e.g., by matching linguistic characteristics). However, distributions that are hard to distinguish on average may still be easy to distinguish for data outliers.
3.2 Blind Attack Techniques
In this section, we introduce simple ‘‘blind’’ attack techniques to distinguish members from non-members in MI evaluations. The datasets we consider consist of either text, or images and text. For simplicity, we focus only on the text modality in both cases. We do not aim for our blind attacks to be optimal. We prioritize simple and interpretable methods to show that existing evaluation datasets suffer from large distribution shifts. Our blind attacks typically aim to achieve a high TPR at a low FPR (i.e., very confident predictions of membership for as many samples as possible).
Date detection.
Some text samples (e.g., from Wikipedia, arXiv, etc.) might contain specific dates. Heuristically, it is unlikely that a text contains specific dates referencing the future. Thus, to place an upper-bound on the date at which a text was written, we simply extract all dates present in the text (using simple regular expressions). We then predict that a sample is a member if all referenced dates fall before some threshold. Such an approach can have false positives when a text sample only references dates far in the past.
Bag-of-words classification.
Explicit dates are just one textual feature we can use. More generally, we can aim to predict membership from arbitrary words in the text. To this end, we train a simple bag-of-words classifier. Here we have to guard against overfitting, since we can always find some word combinations that appear in members and not in non-members (e.g., the exact text of the member samples). So we train our classifier on 80% of the members and non-members, and then evaluate the blind attack on the remaining 20%. As some of the evaluation datasets are small, we aggregate results over a 10-fold cross-validation.
Greedy rare word selection.
The above classifier should work well on average, but might not be optimal at low FPRs. Here, we take a simple greedy approach: we extract all n-grams (for $n\in[1,5]$) and sort them by their TPR-to-FPR ratio on part of the data (i.e., for each n-gram, we compute the fraction of members and non-members that contain this n-gram). We then pick the n-gram with the best ratio, and repeat this procedure. Given a set of selected n-grams, we evaluate our attack on a held-out set by predicting that a sample is a member if it contains any of these n-grams.
4 Case Studies
We now study 8 membership inference datasets proposed for large language models and diffusion models. Table 2 summarizes our results: for each dataset, we create blind attacks that outperform the best existing MI attacks that have access to a trained model. We average the results from multiple runs and report for the same metric used in prior work, either AUC ROC or TPR at low FPR. Whenever possible, we use the exact datasets released by the authors to ensure that no biases are introduced. For datasets that are not publicly available (arXiv-1 month and Gutenberg Meeus et al. (2023)), we follow the specific collection steps outlined in the respective work to create similar datasets.
Таблица 2: Extended version of Table 1, with the results from our blind attacks compared to the best reported MI attack on each evaluation set. We report the same metrics as used in prior work, and add TPR at 1% FPR if no prior results are reported.
Section MI Dataset Metric Best Attack  (%) Ours  (%)
Temporal shifts
 4.1.1 WikiMIA TPR@5%FPR Zhang et al. (2024a) 43.2 Bag of Words 94.7
AUC ROC Zhang et al. (2024a) 83.9 Bag of Words 99.0
4.1.2 BookMIA TPR@5%FPR Zhang et al. (2024b) 33.6 Bag of Words 64.5
AUC ROC Shi et al. (2023) 88.0 Bag of Words 91.4
4.1.3 Temporal Wiki TPR@1%FPR - - Greedy 36.5
AUC ROC Duan et al. (2024) 79.6 Greedy 79.9
4.1.3 Temporal Arxiv TPR@1%FPR - - Bag of Words 9.1
AUC ROC Duan et al. (2024) 74.5 Bag of Words 75.3
4.1.4 Arxiv TPR@1%FPR Meeus et al. (2023) 5.9 Date Detection 10.6
(all vs 1 mo) AUC ROC Meeus et al. (2023) 67.8 Date Detection 72.3
4.1.5 Arxiv TPR@1%FPR Meeus et al. (2024a) 2.5 Greedy 2.7
(1 mo vs 1 mo)
Biased replication
4.2.1 LAION-MI TPR@1%FPR Dubiński et al. (2024) 2.5 Greedy 8.9
4.2.2 Gutenberg TPR@1%FPR Meeus et al. (2023) 18.8 Greedy 55.1
AUC ROC Meeus et al. (2023) 85.6 Bag of Words 96.1
4.1 Temporal Shifts
4.1.1 WikiMIA
The dataset and evaluation.
The WikiMIA dataset Shi et al. (2023) selects members from Wikipedia event pages from before 01/01/2017 and non-members from after 01/01/2023. It thus serves as an MI evaluation set for any LLM trained in between those dates. The best reported prior MI attack on this evaluation set is from the Min-K%++ method of Zhang et al. (2024a).
Our attack.
We first apply our naive date detection attack. We extract all dates in the snippet and check if the latest one is from before 2023. Using this, we obtain a $52.3$% TPR at $5$% FPR. This already beats the state-of-the-art MI attack evaluated on this dataset (see Table 2). To obtain an even stronger blind attack, we train a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. This classifier achieves a near perfect TPR of $94.7$%. (See Appendix 1 for a visualization of the clear distribution shift in the WikiMIA dataset).
4.1.2 BookMIA
The dataset and evaluation.
This dataset Shi et al. (2023) is constructed from 512-token length text snippets from various books. Members are selected from books in the Books3 corpus that have been shown to be memorized by GPT-3. Non-members are taken from books that were first published in or after 2023. In the evaluation of Shi et al. (2023), their Min-Kmethod obtains the highest AUC against GPT-3.
Duarte et al. (2024) extended this dataset to propose the BookTection dataset. Based on their construction, we expected this dataset to suffer from the same drawbacks as BookMIA and thus do not include it in our study.
Our attack.
We train a a bag-of-words classifier on 80% of the dataset, and evaluate on the remaining 20%. As shown in Table 2, a bag-of-words classifier achieves a $91.4$% AUC ROC and beats state-of-the-art MI attacks against GPT-3.111To control for book- and author-specific features (such as character names), we repeat the experiment with no author appearing in both our classifier’s train and test sets. This still results in a significantly above chance AUC ROC of $80.3$%.
4.1.3 Temporal Wiki & arXiv
The datasets and evaluations.
Duan et al. (2024) hypothesize that some MI attacks evaluated on Wiki-MIA may inadvertently exploit temporal shifts between members and non-members. To showcase this, they create datasets that sample members from the Wikipedia and arXiv snippets in the Pile Gao et al. (2020) and non-members from the same sources at a later date. [^2] Duarte et al. (2024) propose a similar ArxivTection dataset constructed using the same temporal principle.
In more detail, in the Temporal arXiv Dataset, members are snippets from arXiv papers posted prior to July 2020 and contained in the Pile training data, while the non-members are sampled from arXiv papers from successively different time ranges after the Pile cutoff of July 2020. The Temporal Wikipedia dataset is also constructed based on the same principle as the WikiMIA dataset, but selects non-member articles from RealTimeData WikiText data created after August 2023 while members are from The Pile from before 03/01/2020. The authors show that existing MI attacks applied to Pythia models improve as the temporal shift increases, with the MI attack of Carlini et al. (2022) performing best.
Our attacks.
We show that even if existing attacks do rely on some temporal features in the Temporal Wikipedia and Temporal arXiv datasets, they do so sub-optimally. We train a bag-of-words classifier which outperforms prior MI results on these datasets. Specifically, we focus on the ‘‘2020-08’’ split (where non-members are taken from arXiv articles in August 2020, right after the Pile cutoff). There, our blind attack slightly outperforms the AUC and TPR at 1% FPR of the best reported MI attack on both datasets (see Table 2).
4.1.4 ArXiv (all vs one month)
The dataset and evaluation.
Meeus et al. (2023) also note that naively re-collecting data from arXiv causes a large temporal shift. Instead, they thus pick the non-members as close after the model’s cutoff date as possible. They build an MI evaluation set by taking all arXiv articles from the RedPajama dataset Together (2023) as members, and all articles from March 2023 (right after the dataset’s cutoff date) as non-members. In contrast to the Temporal arXiv dataset we looked at previously, this dataset uses full arXiv articles rather than just snippets. The authors propose a new MI attack that relies on textual feature extractors, and evaluate it on the OpenLLaMA model Geng & Liu (2023) which was trained on RedPajama.
Our attack.
[…]
4.1.5 ArXiv (one month vs one month)
The dataset and evaluation.
In concurrent work, Meeus et al. (2024a) also note the issue of a temporal shift in their previous dataset we discussed above. To remedy this, they build a new MI evaluation set where the non-members are still arXiv articles from March 2023, but the members are now limited to articles from February 2023 (i.e., right before the cutoff date of the RedPajama dataset Together (2023)). Unfortunately, this dataset is hence suitable for evaluating MI attacks on a very narrow subset of data. The hope is that this aggressive filtering can remove the temporal shift altogether. And indeed, existing MI attacks fare poorly on this split of members and non-members: the authors show that their best MI attack achieves a TPR of $2.5\%$ at an FPR of $1\%$.
Our attack.
Even a one month gap could result in a rather significant distribution shift if some research topics are emerging or disappearing from the mainstream. As an illustrative example, if we wanted to show non-membership, some topics provide a very strong signal: we find that articles from March 2023 are about 25×25$\times$25 × more likely to contain the terms ‘‘GPT-4’’ or ‘‘iccvfinalcopy’’ than articles from February 2023.333GPT-4 was released on March 14th 2023. The submission deadline for ICCV 2023 was on March 8th 2023.
Biases in the opposite direction are not as clear-cut, but still exist. Using $80\%$ of the data, we greedily select individual words that have the highest TPR-to-FPR ratio on this data (top examples include ‘‘pre-registered’’, ‘‘speaker-dependent’’ or ‘‘MMoE’’). We then predict that a sample from the held-out set is a member if and only if it contains one of these words. This yields a blind TPR of $2.7\%$ at an FPR of $1\%$, slightly better than the best MI attack that has access to a trained model. While this TPR is small, it does show that temporal biases are persistent even in simple word choices.
4.2 Biases in Data Replication and Distinguishable Tails
4.2.1 LAION-MI
The dataset.
Dubiński et al. (2024) explicitly try to control for and minimize biases in data replication. However, their techniques still leave distinguishable tails. To create an evaluation dataset for a model trained on LAION Schuhmann et al. (2022), they rely on the multilingual-LAION dataset Schuhmann et al. (2022) to sample non-members. This dataset was sampled from the Web in a similar way to LAION, except that the captions are not in English and the images were not filtered for quality.
To align the two distributions, the authors first apply the same selection process that was used to select images in LAION: (1) they discard all images from multilingual LAION with an ‘‘aesthetics score’’ below some threshold; (2) they translate the non-member captions to English, using Facebook’s M2M100 1.2B. They then train a distinguisher on the image captions, and select a subset of the data where members and non-members are hard to distinguish on average. The authors use a PCA visualization to confirm that the two distributions closely match.
The authors evaluate multiple MI attacks in white-box, grey-box and black-box settings, obtaining a best score of $2.5$% TPR@1%FPR on their LAION-MI dataset.
Our attack.
We posit that outlier members would still be easy to distinguish from non-members. To provide some intuition, we started by looking for individual characters that appear much more frequently in members than in non-members. We show the five characters with the highest distinguishing power in Table 3.
[…]
Таблица 3: Characters with the highest distinguishing power among members and non-members for LAION-MI.
Character Members (%) Non-members (%)
| 0.60 0.00
[…]
0.21 0.00
о 0.23 0.03
т 0.15 0.03
To boost the power of this approach, we apply our greedy selection algorithms to all n-grams of up to 5 characters. We use $80\%$ of the data to pick the n-grams with the highest TPR-to-FPR ratio until we hit a $1\%$ FPR. At that point, we obtain an $8.9$% TPR on the evaluation set—surpassing the best MI attack.
4.2.2 Project Gutenberg
The dataset and evaluation.
This dataset Meeus et al. (2023) consists of books from Project Gutenberg. The members are books contained in the RedPajama dataset Together (2023), more specifically the set of ‘‘PG-19’’ books collected by Rae et al. (2019) which were all originally published before 1919. The non-members consist of books added to the Project Gutenberg repository after the creation of the PG-19 corpus in February 2019. To mitigate the obvious distribution shift in the publication years of members and non-members, Meeus et al. (2023) filter both the members and non-members to only contain books published between 1850 and 1910 (i.e., members are books published between 1850 and 1910, which were added to Project Gutenberg before February 2019. Non-members are books that were originally published in the same period, but added to Project Gutenberg after February 2019). The authors evaluate a new MI attack (see Section 4.1.4 for details) against the OpenLLaMA model Geng & Liu (2023) trained on RedPajama.
Our attack.
While the publication dates of the members and non-members are similarly distributed, we find that the date on which a book is added to Project Gutenberg still introduces a noticeable distribution shift. Indeed, it appears that the format of the preface metadata that Project Gutenberg adds to books has changed at some point between the collection of PG-19 and the non-member corpus (e.g., the link to the HTML version of a book used to have a .htm file extension, and this was changed to a .html extension for books added after a certain date). By exploiting such small discrepancies, our greedy n-gram attack applied to the first 1,000 characters of a book achieves $55.1\%$ TPR at a $1\%$ FPR, 3 times higher than the best MI attack evaluated on this dataset.
Of course, it may be possible to preprocess this particular dataset to get rid of these formatting discrepancies. Nevertheless, we believe our results illustrate how brittle the process of re-collecting an identically distributed non-member set can be. Moreover, even if we ignore any formatting information, and apply our attack directly to the book text, we can still distinguish members and non-members far better than chance, with a TPR of $16.6\%$ at a $1\%$ FPR.
5 A Path Forward: the Pile, DataComp and DataComp-LM
As we have seen, building MI evaluation datasets a posteriori is incredibly challenging, due to the many possibilities for distribution shift. It may be possible to apply more rigorous filtering techniques to minimize these shifts, so that even strong blind attacks cannot reliably distinguish members and non-members. Any MI attack’s performance should then be compared against these strong blind baselines.
An alternative avenue is to forego MI evaluations on arbitrary foundation models, and focus on those models for which identically distributed sets of members and non-members are available. A prime example is the Pile Gao et al. (2020), which has an official test set. Models trained on the Pile training set—e.g., Pythia Biderman et al. (2023) or GPT-NeoX-20B Black et al. (2022)—are thus a popular target for MI attack evaluations (see e.g., Duan et al. (2024); Maini et al. (2024); Li et al. (2023)). Unfortunately, the Pile only contains text, and is too small to train state-of-the-art language models.
A more recent and broadly applicable alternative is to use models trained on the DataComp Gadre et al. (2024) and DataComp-LM Li et al. (2024) benchmarks. These benchmarks contribute multiple models (ViTs and LLMs) trained on data sampled randomly from a larger dataset. More precisely, these benchmarks introduce two datasets, CommonPool and DCLM-POOL, which contain respectively 12.8 billion image-text pairs and 240 trillion text tokens sampled from Common Crawl. To enable experiments at smaller budgets, the authors produce smaller data pools sampled randomly from the full pool. They then train a variety of models on each pool (see Table 4). [^4] https://github.com/mlfoundations/datacomp, [^5] https://github.com/mlfoundations/dclm.
A minor complication is that most models in the DataComp and DataComp-LM benchmarks are not trained on the entirety of a pool. Rather, the benchmarks encourage models to be trained on filtered datasets: given a pool of data $\mathcal{P}$ and a binary filter $f:\mathcal{X}\mapsto\{0,1\}$, models are trained on the filtered pool $\{x\in\mathcal{P}\mid f(x)\}$. This training set is thus no longer a random subset of the full dataset. But this is easily remedied by applying exactly the same filter $f$ to select the non-members.
We encourage future work on MI attacks for foundation models to evaluate their methods on these models and datasets.
Таблица 4: Models trained on datasets with random train-test splits for DataComp Gadre et al. (2024) and DataComp-LM Li et al. (2024). For each dataset pool, we report the pool’s absolute size (in image-text pairs, respectively tokens) and its relative size compared to the global pool that it was randomly sampled from. Gadre et al. (2024) and Li et al. (2024) have pretrained ViTs and LLMs on each pool size.
Modality Pool Name Pool Size Models
Vision+Text CommonPool Large 1.280B. (10%) ViT-B/16
CommonPool Medium 0.128B0. (1%) ViT-B/32
CommonPool Small 0.013B (0.1%) ViT-B/32
Text DCLM 400M-1x 0.47T (0.2%) 412M LLM
DCLM 1B-1x 1.64T (0.2%) 1.4B LLM
DCLM 1B-5x 8.20T (3.4%) 1.4B LLM
DCLM 7B-1x 7.85T (3.3%) 6.9B LLM
DCLM 7B-5x 15.7T (6.5%) 6.9B LLM
As part of their experiments, Gadre et al. (2024) and Li et al. (2024) trained CLIP models of various sizes and a 1B parameter language model on random subsets of these datasets. These models and datasets might thus form an ideal testbed for MI evaluations on web-scale text and vision models.
Future web-scale dataset and model developers could similarly release an official IID (independent and identically distributed) split to enable rigorous MI evaluations for a wider variety of models and modalities.
6 Conclusion
We have shown that current evaluations of MI attacks for foundation models cannot be trusted, as members and non-members can be reliably distinguished by simple blind attacks with no knowledge of the model. State-of-the-art MI attacks may thus be ineffective at extracting any actual membership information, and cannot be relied on for applications such as copyright detection or auditing unlearning methods. The MI evaluation datasets we investigated should likely be discarded due to the large distribution shifts between members and non-members, and replaced by datasets with a random train-test split—e.g., the Pile or DataComp.
7 Reproducibility
Our results can be reproduced using our code-base available at: https://github.com/ethz-spylab/Blind-MIA. The code-base includes most of the datasets and scripts to download the datasets that were too large to include.
Список литературы
  • Biderman et al. (2023) Stella Biderman, Hailey Schoelkopf, Quentin Gregory Anthony, Herbie Bradley, Kyle O’Brien, Eric Hallahan, Mohammad Aflah Khan, Shivanshu Purohit, USVSN Sai Prashanth, Edward Raff, et al. Pythia: A suite for analyzing large language models across training and scaling. In International Conference on Machine Learning, pp.  2397–2430. PMLR, 2023.
  • Black et al. (2022) Sid Black, Stella Biderman, Eric Hallahan, Quentin Anthony, Leo Gao, Laurence Golding, Horace He, Connor Leahy, Kyle McDonell, Jason Phang, et al. Gpt-neox-20b: An open-source autoregressive language model. arXiv preprint arXiv:2204.06745, 2022.
  • Bommasani et al. (2021) Rishi Bommasani, Drew A Hudson, Ehsan Adeli, Russ Altman, Simran Arora, Sydney von Arx, Michael S Bernstein, Jeannette Bohg, Antoine Bosselut, Emma Brunskill, et al. On the opportunities and risks of foundation models. arXiv preprint arXiv:2108.07258, 2021.
  • Bourtoule et al. (2021) Lucas Bourtoule, Varun Chandrasekaran, Christopher A Choquette-Choo, Hengrui Jia, Adelin Travers, Baiwu Zhang, David Lie, and Nicolas Papernot. Machine unlearning. In 2021 IEEE Symposium on Security and Privacy (SP), pp.  141–159. IEEE, 2021.
  • Carlini et al. (2021) Nicholas Carlini, Florian Tramer, Eric Wallace, Matthew Jagielski, Ariel Herbert-Voss, Katherine Lee, Adam Roberts, Tom Brown, Dawn Song, Ulfar Erlingsson, et al. Extracting training data from large language models. In 30th USENIX Security Symposium (USENIX Security 21), pp.  2633–2650, 2021.
  • Carlini et al. (2022) Nicholas Carlini, Steve Chien, Milad Nasr, Shuang Song, Andreas Terzis, and Florian Tramer. Membership inference attacks from first principles. In 2022 IEEE Symposium on Security and Privacy (SP), pp.  1897–1914. IEEE, 2022.
  • Duan et al. (2024) Michael Duan, Anshuman Suri, Niloofar Mireshghallah, Sewon Min, Weijia Shi, Luke Zettlemoyer, Yulia Tsvetkov, Yejin Choi, David Evans, and Hannaneh Hajishirzi. Do membership inference attacks work on large language models? arXiv preprint arXiv:2402.07841, 2024.
  • Duarte et al. (2024) André V Duarte, Xuandong Zhao, Arlindo L Oliveira, and Lei Li. De-cop: Detecting copyrighted content in language models training data. arXiv preprint arXiv:2402.09910, 2024.
  • Dubiński et al. (2024) Jan Dubiński, Antoni Kowalczuk, Stanisław Pawlak, Przemyslaw Rokita, Tomasz Trzciński, and Paweł Morawiecki. Towards more realistic membership inference attacks on large diffusion models. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision, pp.  4860–4869, 2024.
  • Engstrom et al. (2020) Logan Engstrom, Andrew Ilyas, Shibani Santurkar, Dimitris Tsipras, Jacob Steinhardt, and Aleksander Madry. Identifying statistical bias in dataset replication. In International Conference on Machine Learning, pp.  2922–2932. PMLR, 2020.
  • Gadre et al. (2024) Samir Yitzhak Gadre, Gabriel Ilharco, Alex Fang, Jonathan Hayase, Georgios Smyrnis, Thao Nguyen, Ryan Marten, Mitchell Wortsman, Dhruba Ghosh, Jieyu Zhang, et al. DataComp: In search of the next generation of multimodal datasets. Advances in Neural Information Processing Systems, 36, 2024.
  • Gao et al. (2020) Leo Gao, Stella Biderman, Sid Black, Laurence Golding, Travis Hoppe, Charles Foster, Jason Phang, Horace He, Anish Thite, Noa Nabeshima, et al. The Pile: An 800GB dataset of diverse text for language modeling. arXiv preprint arXiv:2101.00027, 2020.
  • Gemini Team (2023) Gemini Team. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805, 2023.
  • Geng & Liu (2023) Xinyang Geng and Hao Liu. OpenLLaMA: An open reproduction of LLaMA, May 2023. URL https://github.com/openlm-research/open_llama.
  • Jayaraman et al. (2020) Bargav Jayaraman, Lingxiao Wang, Katherine Knipmeyer, Quanquan Gu, and David Evans. Revisiting membership inference under realistic assumptions. arXiv preprint arXiv:2005.10881, 2020.
  • Ko et al. (2023) Myeongseob Ko, Ming Jin, Chenguang Wang, and Ruoxi Jia. Practical membership inference attacks against large-scale multi-modal models: A pilot study. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pp.  4871–4881, 2023.
  • Li et al. (2024) Jeffrey Li, Alex Fang, Georgios Smyrnis, Maor Ivgi, Matt Jordan, Samir Gadre, Hritik Bansal, Etash Guha, Sedrick Keh, Kushal Arora, et al. DataComp-LM: In search of the next generation of training sets for language models. arXiv preprint arXiv:2406.11794, 2024.
  • Li et al. (2023) Marvin Li, Jason Wang, Jeffrey Wang, and Seth Neel. MoPe: Model perturbation-based privacy attacks on language models. arXiv preprint arXiv:2310.14369, 2023.
  • Maini et al. (2024) Pratyush Maini, Hengrui Jia, Nicolas Papernot, and Adam Dziedzic. LLM dataset inference: Did you train on my dataset?, 2024.
  • Meeus et al. (2023) Matthieu Meeus, Shubham Jain, Marek Rei, and Yves-Alexandre de Montjoye. Did the neurons read your book? Document-level membership inference for large language models. arXiv preprint arXiv:2310.15007, 2023.
  • Meeus et al. (2024a) Matthieu Meeus, Shubham Jain, Marek Rei, and Yves-Alexandre de Montjoye. Inherent challenges of post-hoc membership inference for large language models. arXiv preprint arXiv:2406.17975, 2024a.
  • Meeus et al. (2024b) Matthieu Meeus, Igor Shilov, Manuel Faysse, and Yves-Alexandre de Montjoye. Copyright traps for large language models. arXiv preprint arXiv:2402.09363, 2024b.
  • OpenAI (2023) OpenAI. GPT-4 technical report. arXiv preprint arXiv:2303.08774, 2023.
  • Oren et al. (2023) Yonatan Oren, Nicole Meister, Niladri Chatterji, Faisal Ladhak, and Tatsunori B Hashimoto. Proving test set contamination in black box language models. arXiv preprint arXiv:2310.17623, 2023.
  • Panaitescu-Liess et al. (2024) Michael-Andrei Panaitescu-Liess, Zora Che, Bang An, Yuancheng Xu, Pankayaraj Pathmanathan, Souradip Chakraborty, Sicheng Zhu, Tom Goldstein, and Furong Huang. Can watermarking large language models prevent copyrighted text generation and hide training data? arXiv preprint arXiv:2407.17417, 2024.
  • Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. Language models are unsupervised multitask learners. OpenAI blog, 1(8):9, 2019.
  • Rae et al. (2019) Jack W Rae, Anna Potapenko, Siddhant M Jayakumar, and Timothy P Lillicrap. Compressive transformers for long-range sequence modelling. arXiv preprint arXiv:1911.05507, 2019.
  • Raffel et al. (2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J Liu. Exploring the limits of transfer learning with a unified text-to-text transformer. Journal of machine learning research, 21(140):1–67, 2020.
  • Ramesh et al. (2021) Aditya Ramesh, Mikhail Pavlov, Gabriel Goh, Scott Gray, Chelsea Voss, Alec Radford, Mark Chen, and Ilya Sutskever. Zero-shot text-to-image generation. In International Conference on Machine Learning, 2021.
  • Recht et al. (2019) Benjamin Recht, Rebecca Roelofs, Ludwig Schmidt, and Vaishaal Shankar. Do ImageNet classifiers generalize to ImageNet? In International conference on machine learning, pp.  5389–5400. PMLR, 2019.
  • Schuhmann et al. (2022) Christoph Schuhmann, Romain Beaumont, Richard Vencu, Cade Gordon, Ross Wightman, Mehdi Cherti, Theo Coombes, Aarush Katta, Clayton Mullis, Mitchell Wortsman, et al. LAION-5B: An open large-scale dataset for training next generation image-text models. Advances in Neural Information Processing Systems, 35:25278–25294, 2022.
  • Shi et al. (2023) Weijia Shi, Anirudh Ajith, Mengzhou Xia, Yangsibo Huang, Daogao Liu, Terra Blevins, Danqi Chen, and Luke Zettlemoyer. Detecting pretraining data from large language models. arXiv preprint arXiv:2310.16789, 2023.
  • Shokri et al. (2017) Reza Shokri, Marco Stronati, Congzheng Song, and Vitaly Shmatikov. Membership inference attacks against machine learning models. In 2017 IEEE symposium on security and privacy (SP), pp.  3–18. IEEE, 2017.
  • Soldaini et al. (2024) Luca Soldaini, Rodney Kinney, Akshita Bhagia, Dustin Schwenk, David Atkinson, Russell Authur, Ben Bogin, Khyathi Chandu, Jennifer Dumas, Yanai Elazar, et al. Dolma: An open corpus of three trillion tokens for language model pretraining research. arXiv preprint arXiv:2402.00159, 2024.
  • Together (2023) Together. RedPajama: an open dataset for training large language models, 2023. URL https://github.com/togethercomputer/RedPajama-Data.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. LLaMA: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971, 2023.
  • Yeom et al. (2018) Samuel Yeom, Irene Giacomelli, Matt Fredrikson, and Somesh Jha. Privacy risk in machine learning: Analyzing the connection to overfitting. In 2018 IEEE 31st computer security foundations symposium (CSF), pp.  268–282. IEEE, 2018.
  • Zhang et al. (2024a) Jingyang Zhang, Jingwei Sun, Eric Yeats, Yang Ouyang, Martin Kuo, Jianyi Zhang, Hao Yang, and Hai Li. Min-K%++: Improved baseline for detecting pre-training data from large language models. arXiv preprint arXiv:2404.02936, 2024a.
  • Zhang et al. (2024b) Weichao Zhang, Ruqing Zhang, Jiafeng Guo, Maarten de Rijke, Yixing Fan, and Xueqi Cheng. Pretraining data detection for large language models: A divergence-based calibration method. arXiv preprint arXiv:2409.14781, 2024b.
Приложение A Appendix
Рис. 1: In some cases, the distribution shift is easy to visualise. Here is the PCA plot of the WikiMIA dataset which shows members and non-members forming distinguishable clusters.